# Distilling Implicit Multimodal Knowledge into LLMs for Zero-Resource Dialogue Generation

### v1

@@ Abstract
Integrating multimodal knowledge into large language models (LLMs) represents a significant advancement in dialogue generation capabilities. However, the effective incorporation of such knowledge in zero-resource scenarios remains a substantial challenge due to the scarcity of diverse, high-quality dialogue datasets. To address this, we propose the Visual Implicit Knowledge Distillation Framework (VIKDF), an innovative approach aimed at enhancing LLMs for enriched dialogue generation in zero-resource contexts by leveraging implicit multimodal knowledge. VIKDF comprises two main stages: knowledge distillation, using an Implicit Query Transformer to extract and encode visual implicit knowledge from image-text pairs into knowledge vectors; and knowledge integration, employing a novel Bidirectional Variational Information Fusion technique to seamlessly integrate these distilled vectors into LLMs. This enables the LLMs to generate dialogues that are not only coherent and engaging but also exhibit a deep understanding of the context through implicit multimodal cues, effectively overcoming the limitations of zero-resource scenarios. Our extensive experimentation across two dialogue datasets shows that VIKDF outperforms existing state-of-the-art models in generating high-quality dialogues. The code will be publicly available following acceptance.
Index Terms:
Large language models, Multimodal knowledge, Zero resource, Dialogue generation.
I Introduction
@@
The integration of multimodal knowledge into dialogue systems signifies a major progression towards more nuanced and human-like communication capabilities. It enables these systems to understand and interpret the nuances of human communication that transcend beyond mere text, capturing the essence of multimodal interactions [4, 5, 6]. Building upon this concept, there has been an increase in research focused on augmenting LLMs with multimodal knowledge [7, 8, 9, 10]. This involves processing and understanding information across different modalities, such as images, videos, and audio, thereby equipping LLMs to perform tasks that necessitate cross-modal comprehension. While these developments significantly expand the capabilities of LLMs in engaging with multimodal content, challenges persist in effectively applying multimodal knowledge in dialogue generation, necessitating further exploration and innovation in this area.
@@
II Related Work
II-A Multimodal Dialogue Generation
@@
II-B Multimodal Knowledge Distillation and Integration
@@
Prior research on this topic has predominantly concentrated on explicit multimodal information, such as the direct fusion of image and textual features [5, 26, 27, 28] or employing pre-trained multimodal models [29] to extract multimodal representations [19, 30]. However, these methods face limitations when applied to LLMs. Li et al. [7] proposed an innovative solution, the Q-Former, which uses learnable query vectors to distill multimodal knowledge and integrate it into LLMs. Although this method has shown promising results in improving the multimodal capabilities of LLMs, it still encounters challenges in handling implicit multimodal knowledge, especially for dialogue generation.
@@
III Methodology
III-A Task Formalization and Model Architecture
@@
III-A1 Text Encoder and Image Encoder
@@
III-A2 Implicit Query Transformer
The IQ-Former is a specially designed querying transformer that distills visual implicit knowledge from encoded image-text pairs. As illustrated in Figure 2, the IQ-Former is structured as a transformer encoder equipped with a cross-attention mechanism. Uniquely, its inputs are a set of learnable query vectors $\mathbf{Q}={\mathbf{q}_{1},\mathbf{q}_{2},...,\mathbf{q}_{n}}$, where $n$ is the number of queries. The IQ-Former uses the cross-attention mechanism to interact between the query vectors and both $\mathbf{H}_{T}$ and $\mathbf{H}_{I}$, and generates two sets of knowledge vectors $\mathbf{K}_{T}={\mathbf{k}_{T_{1}},\mathbf{k}_{T_{2}},...,\mathbf{k}_{T_{n}}}$ and $\mathbf{K}_{I}={\mathbf{k}_{I_{1}},\mathbf{k}_{I_{2}},...,\mathbf{k}_{I_{n}}}$, respectively. The vectors in $\mathbf{K}_{T}$ contain enriched textual information, whereas those in $\mathbf{K}_{I}$ are infused with explicit visual information. By optimizing the IQ-Former with specifically designed objectives, as described in Section III-B, it can be efficiently distill visual implicit knowledge into $\mathbf{K}_{T}$.
III-A3 Large Language Model
The LLM is tasked with generating dialogue responses based on the dialogue context and the distilled visual implicit knowledge. We employ the Llama-2 model [33] as our LLM, a pre-trained autoregressive language model renowned for its capability to produce natural and varied text. The LLM takes the dialogue context $C$ and distilled visual implicit knowledge $\mathbf{K}_{T}$ as inputs, and generates the response $R$ as output. To facilitate the integration of $\mathbf{K}_{T}$ into the LLM, we introduce a novel technique termed Bidirectional Variational Information Fusion, detailed further in Section III-C. The LLM’s parameters are also frozen during training.
@@ III-A3 Large Language Model
III-B Knowledge Distillation Stage
@@
III-B1 Text-Image Matching
@@
III-B2 Text-Assisted Masked Image Modeling
@@
III-B3 Image-Assisted Masked Text Modeling
@@
III-C Knowledge Integration Stage
@@
III-C1 Instruction-aware Contextual Inference
@@
III-C2 Instruction-aware Knowledge Reconstruction
@@
III-D Zero-Resource Learning Detail
III-D1 Training
To train our framework, we use four large-scale image-text datasets, including COCO Captions [36], CC3M [37], CC12M [38], and SBU [39]. We follow the same data processing and augmentation methods as BLIP-2 [7], which generates synthetic captions for web images using a pre-trained captioning model and a CLIP model. These datasets contain tens of millions of image-text pairs that cover a wide range of topics and scenarios, providing a rich source of visual implicit knowledge. We concurrently train both stages of our framework, employing image-text pair data in accordance with Equations (5) and (9). For fine-tuning on specific tasks, we minimize the negative log-likelihood loss to optimize the probability $P(R|C,K)$, as detailed in Section III-D2.
III-D2 Inference
To perform zero-resource inference, we leverage the trained IQ-Former and the LLM to produce dialogue responses that enriched with the visual implicit knowledge distilled from the dialogue context. Given a dialogue context $C$, we first pass it through the IQ-Former with the text encoder to model $P(K|C)$, thereby acquiring knowledge vectors $K$ that encapsulate visual implicit knowledge derived from $C$. Next, we input $K$ and $C$ into the LLM, which models $P(R|C,K)$, and then generate the response $R$ by sampling from the probability distribution over the vocabulary. Since the LLM has learned to integrate the visual implicit knowledge into the dialogue generation, it can produce natural and engaging responses that are consistent with the visual context, even in the absence of any explicit multimodal inputs.
IV Experiments
@@ IV-A Datasets
Similar to [14], we use the text-based dialogue dataset Reddit Conversation to train the model’s foundational dialogue capabilities, and the image-grounded dialogue dataset Image-Chat to validate the model’s zero-resource dialogue generation capabilities.
IV-B Implementation Details
@@
VIKDF adopts a simultaneous training regimen for both knowledge distillation and integration stages. This strategy is executed on four NVIDIA RTX 4090 GPUs, utilizing a batch size of 128 across 100,000 training steps. For optimization, we employ mixed-precision training with bfloat16 and utilize the AdamW optimizer [40] with a learning rate of $1e^{-4}$. The learning rate is accompanied by a weight decay of 0.05 and a linear learning rate warmup over the first $10\%$ of the total steps. The hyperparameters $\lambda_{1}$ through $\lambda_{5}$ are set as 0.5, 0.2, 0.2, 1, and 0.5, respectively. For the set of learnable query vectors within the IQ-Former, we set $n$ to 32, with each vector having a dimensionality of 768. In the text-assisted masked image modeling, we apply a random masking strategy with $28\times 28$ patches at a 0.6 ratio. For the image-assisted masked text modeling, the strategy involves a 0.15 mask ratio.
IV-C Baseline Models
@@
  • BART [42], a pre-eminent sequence-to-sequence pre-training model that utilizes a Transformer architecture.
@@
Among them, ImgVAE, Maria, and ZRIGF are multimodal models, whereas the rest are unimodal text-based models. We prompt Llama-2 following the approach outlined in the Hugging Face blog [^1] https://huggingface.co/blog/llama2, and employ ChatGPT in accordance with the methodology described in [14].
IV-D Evaluation Metrics
@@
Task Methods PPL BLEU-1 ROUGE-L Average Extrema Greedy      Dis-1 Dis-2
Reddit Conversation Seq2Seq $77.27$ $12.21$ $10.81$ $78.38$ $40.06$ $62.64$ $0.53$ $1.96$
BART $44.73$ $13.51$ $12.50$ $80.21$ $41.63$ $63.72$ $4.17$ $16.98$
Llama-2 $155.69$ $10.27$ $10.52$ $81.72$ $35.95$ $60.70$ $4.94$ $31.19$
ChatGPT - $11.62$ $11.29$ 82.39 $37.48$ $62.05$ $5.28$ 38.63
ImgVAE $72.06$ $12.58$ $12.05$ $79.95$ $42.38$ $63.55$ $1.52$ $6.34$
Maria $56.23$ $14.10$ $12.66$ $81.76$ $43.04$ $63.98$ $4.83$ $22.87$
ZRIGF 36.21 16.06 14.51 $82.27$ 43.79 64.53     5.79 $26.57$
VIKDF 15.01 16.41 14.82 82.53 43.71 64.88     6.39 35.84
Image-Chat Seq2Seq $50.82$ $11.34$ $13.65$ $82.95$ $47.45$ $65.67$ $1.28$ $7.80$
BART $37.26$ $13.41$ $14.24$ $84.48$ $48.57$ $66.49$ $2.44$ $15.79$
Llama-2 $193.20$ $9.93$ $11.56$ $85.44$ $40.18$ $63.05$ $4.69$ $30.81$
ChatGPT - $10.77$ $11.62$ $86.17$ $43.02$ $64.66$ $5.32$ 37.77
ImgVAE $41.94$ $16.07$ $15.98$ $85.81$ $49.59$ $67.44$ $1.68$ $7.22$
Maria $37.49$ $14.74$ $14.59$ $85.72$ $50.58$ $66.89$ $2.57$ $11.99$
@@
ZRIGF $29.82$ $16.86$ 17.21 86.30 51.41 68.56 $2.59$ $10.62$
[…]
VIKDF 12.56 17.92 17.33 86.47 50.83 68.45 $4.61$ $23.30$
1/4 Data 12.84 17.81 $16.91$ $86.07$ $50.01$ $67.55$ $4.40$ $21.53$
1/8 Data $13.12$ $17.70$ $16.84$ $85.98$ $49.76$ $67.35$ $4.28$ $20.54$
Zero Data $27.32$ $15.35$ $15.30$ $85.04$ $45.86$ $66.02$     6.17 32.93
V Results and Discussion
[…]
V-A Automatic Evaluation Results
Our proposed VIKDF demonstrates outstanding performance in zero-resource dialogue generation, outperforming both traditional and multimodal baseline models across various metrics on the Reddit Conversation and Image-Chat datasets. Table I presents a comprehensive comparison of the automated evaluation metrics.
In the Reddit Conversation task, VIKDF achieves a significantly lower perplexity of 15.01, indicating superior fluency in generated dialogues compared to the nearest competitor, ZRIGF, which records a PPL of 36.21 in scenarios with explicit image guidance. Moreover, VIKDF outperforms all baselines in BLEU-1 and ROUGE-L scores, achieving 16.41 and 14.82, respectively, which highlights its ability to generate responses that are closely aligned with human references. Additionally, VIKDF surpasses other models in both Average and Greedy semantic similarity metrics, underscoring its enhanced ability to maintain semantic coherence in dialogues. Despite scoring slightly lower on the Extrema metric compared to ZRIGF, VIKDF remains competitive. Although it has a marginally lower Dis-2 score than ChatGPT, VIKDF’s strong performance in generating diverse dialogues is evident. This suggests that while maintaining high relevance and accuracy, VIKDF also generates a wide range of responses, contributing to more dynamic and engaging dialogues. The outstanding performance of VIKDF in the Reddit Conversation task highlights the substantial benefits of incorporating implicit knowledge into purely text-based dialogue systems, proving that a deep fusion of visual and contextual understanding significantly enhances the quality and engagement of the dialogues generated.
[…]
TABLE II: Human evaluation outcomes for the Image-Chat dataset in a fully zero-resource scenario.
@@ V-A Automatic Evaluation Results
V-B Human Evaluation Results

### v2

# Distilling Implicit Multimodal Knowledge into Large Language Models for Zero-Resource Dialogue Generation

@@ Abstract
Integrating multimodal knowledge into large language models (LLMs) represents a significant advancement in dialogue generation capabilities. However, the effective incorporation of such knowledge in zero-resource scenarios remains a substantial challenge due to the scarcity of diverse, high-quality dialogue datasets. To address this, we propose the Visual Implicit Knowledge Distillation Framework (VIKDF), an innovative approach aimed at enhancing LLMs for enriched dialogue generation in zero-resource contexts by leveraging implicit multimodal knowledge. VIKDF comprises two main stages: knowledge distillation, using an Implicit Query Transformer to extract and encode visual implicit knowledge from image-text pairs into knowledge vectors; and knowledge integration, employing a novel Bidirectional Variational Information Fusion technique to seamlessly integrate these distilled vectors into LLMs. This enables the LLMs to generate dialogues that are not only coherent and engaging but also exhibit a deep understanding of the context through implicit multimodal cues, effectively overcoming the limitations of zero-resource scenarios. Our extensive experimentation across two dialogue datasets shows that VIKDF outperforms existing state-of-the-art models in generating high-quality dialogues. The code is available at https://github.com/zhangbo-nlp/VIKDF.
keywords:
Large language models , Multimodal fusion , Zero resource , Dialogue generation
journal: Information Fusion
[…]
[label1]organization=School of Computer Science and Technology, Dalian University of Technology,city=Dalian, postcode=116024, state=Liaoning, country=China
[…]
[label2]organization=School of Computer Science and Information Engineering, Hefei University of Technology,city=Hefei, postcode=230601, state=Anhui, country=China
1 Introduction
@@
The integration of multimodal knowledge into dialogue systems signifies a major progression towards more nuanced and human-like communication capabilities. It enables these systems to understand and interpret the nuances of human communication that transcend beyond mere text, capturing the essence of multimodal interactions [4, 5, 6, 7]. Building upon this concept, there has been an increase in research focused on augmenting LLMs with multimodal knowledge [8, 9, 10, 11]. This involves processing and understanding information across different modalities, such as images, videos, and audio, thereby equipping LLMs to perform tasks that necessitate cross-modal comprehension. While these developments significantly expand the capabilities of LLMs in engaging with multimodal content, challenges persist in effectively applying multimodal knowledge in dialogue generation, necessitating further exploration and innovation in this area.
@@
2 Related Work
2.1 Multimodal Dialogue Generation
@@
2.2 Multimodal Knowledge Distillation and Integration
@@
Prior research on this topic has predominantly concentrated on explicit multimodal information, such as the direct fusion of image and textual features [5, 27, 28, 29] or employing pre-trained multimodal models [30] to extract multimodal representations [20, 31]. However, these methods face limitations when applied to LLMs. Vision-Language Models (VLMs), like LLaVA [32] and Mini-Gemini [33], are designed to process explicit multimodal inputs by interpreting both visual and textual modalities together. They typically train on large multimodal datasets to maximize alignment and shared understanding between vision and language components, enabling comprehensive tasks like visual question answering, image captioning, and scene understanding. Moreover, Li et al. [8] proposed an innovative solution, the Q-Former, which uses learnable query vectors to distill multimodal knowledge and integrate it into LLMs. Although this method has shown promising results in improving the multimodal capabilities of LLMs, it still encounters challenges in handling implicit multimodal knowledge, especially for dialogue generation.
@@
3 Methodology
3.1 Task Formalization and Model Architecture
@@
3.1.1 Text Encoder and Image Encoder
@@
3.1.2 Implicit Query Transformer
The IQ-Former is a specially designed querying transformer that distills visual implicit knowledge from encoded image-text pairs. As illustrated in Figure 2, the IQ-Former is structured as a transformer encoder equipped with a cross-attention mechanism. Uniquely, its inputs are a set of learnable query vectors $\mathbf{Q}={\mathbf{q}_{1},\mathbf{q}_{2},...,\mathbf{q}_{n}}$, where $n$ is the number of queries. The IQ-Former uses the cross-attention mechanism to interact between the query vectors and both $\mathbf{H}_{T}$ and $\mathbf{H}_{I}$, and generates two sets of knowledge vectors $\mathbf{K}_{T}={\mathbf{k}_{T_{1}},\mathbf{k}_{T_{2}},...,\mathbf{k}_{T_{n}}}$ and $\mathbf{K}_{I}={\mathbf{k}_{I_{1}},\mathbf{k}_{I_{2}},...,\mathbf{k}_{I_{n}}}$, respectively. The vectors in $\mathbf{K}_{T}$ contain enriched textual information, whereas those in $\mathbf{K}_{I}$ are infused with explicit visual information. By optimizing the IQ-Former with specifically designed objectives, as described in Section 3.2, it can be efficiently distill visual implicit knowledge into $\mathbf{K}_{T}$.
3.1.3 Large Language Model
The LLM is tasked with generating dialogue responses based on the dialogue context and the distilled visual implicit knowledge. We employ the Llama-2 model [36] as our LLM, a pre-trained autoregressive language model renowned for its capability to produce natural and varied text. The LLM takes the dialogue context $C$ and distilled visual implicit knowledge $\mathbf{K}_{T}$ as inputs, and generates the response $R$ as output. To facilitate the integration of $\mathbf{K}_{T}$ into the LLM, we introduce a novel technique termed Bidirectional Variational Information Fusion, detailed further in Section 3.3. The LLM’s parameters are also frozen during training.
@@ 3.1.3 Large Language Model
3.2 Knowledge Distillation Stage
@@
3.2.1 Text-Image Matching
@@
3.2.2 Text-Assisted Masked Image Modeling
@@
3.2.3 Image-Assisted Masked Text Modeling
@@
3.3 Knowledge Integration Stage
@@
3.3.1 Instruction-aware Contextual Inference
@@
3.3.2 Instruction-aware Knowledge Reconstruction
@@
3.4 Zero-Resource Learning Detail
3.4.1 Training
To train our framework, we initially use four large-scale image-text datasets, including COCO Captions [39], CC3M [40], CC12M [41], and SBU [42], for pre-training the IQ-Former and aligning it with the LLM. We follow the same data processing and augmentation methods as BLIP-2 [8], which generates synthetic captions for web images using a pre-trained captioning model and a CLIP model. These datasets contain tens of millions of image-text pairs that cover a wide range of topics and scenarios, providing a rich source of visual implicit knowledge. We concurrently train both stages of our framework, employing image-text pair data in accordance with Equations (5) and (9).
For fine-tuning on specific dialogue tasks, we utilize only textual data without any image input, aligning with our goal of zero-resource scenarios where multimodal data is unavailable during fine-tuning and inference. During this phase, the IQ-Former distills visual implicit knowledge $K$ from the dialogue context $C$ alone by modeling $P(K|C)$. The fine-tuning process optimizes the negative log-likelihood loss to enhance the probability $P(R|C,K)$, as detailed in Section 3.4.2.
3.4.2 Inference
To perform zero-resource inference, we leverage the trained IQ-Former and the LLM to produce dialogue responses enriched with the visual implicit knowledge distilled from the dialogue context. Given a dialogue context $C$, we first pass it through the IQ-Former with the text encoder to model $P(K|C)$, thereby acquiring knowledge vectors $K$ that encapsulate visual implicit knowledge derived from $C$. Notably, no image input is provided at this stage; the IQ-Former operates solely on the textual input $C$. Next, we input $K$ and $C$ into the LLM, which models $P(R|C,K)$, and then generate the response $R$ by sampling from the probability distribution over the vocabulary. Since the LLM has learned to integrate the visual implicit knowledge into the dialogue generation during training, it can produce natural and engaging responses that are consistent with the visual context implied by $C$, even in the absence of any explicit multimodal inputs.
4 Experiments
@@ 4.1 Datasets
Similar to [15], we fine-tune the model using the Reddit Conversation dataset to enhance the foundational dialogue generation capabilities of the LLM. To evaluate the model’s zero-resource dialogue generation capabilities, we directly perform inference on the Image-Chat dataset after fine-tuning on Reddit Conversation. Additionally, we conduct experiments by fine-tuning the model on Image-Chat without providing any images.
4.2 Implementation Details
@@
VIKDF adopts a simultaneous training regimen for both knowledge distillation and integration stages. This strategy is executed on four NVIDIA RTX 4090 GPUs, utilizing a batch size of 128 across 100,000 training steps. For optimization, we employ mixed-precision training with bfloat16 and utilize the AdamW optimizer [43] with a learning rate of $1e^{-4}$. The learning rate is accompanied by a weight decay of 0.05 and a linear learning rate warmup over the first $10\%$ of the total steps. The hyperparameters $\lambda_{1}$ through $\lambda_{5}$ are set as 0.5, 0.2, 0.2, 1, and 0.5, respectively. For the set of query vectors within the IQ-Former, we set $n$ to 32, with each vector having a dimensionality of 768. In the text-assisted masked image modeling, we apply a random masking strategy with $28\times 28$ patches at a 0.6 ratio. For the image-assisted masked text modeling, the strategy involves a 0.15 mask ratio.
[…]
Task Methods #Tuned Params PPL BLEU-1 ROUGE-L Average Extrema Greedy Dis-1 Dis-2
Reddit Conversation Seq2Seq 24M 77.27 12.21 10.81 78.38 40.06 62.64 0.53 1.96
BART 406M 44.73 13.51 12.50 80.21 41.63 63.72 4.17 16.98
Llama-2 0M 155.69 10.27 10.52 81.72 35.95 60.70 4.94 31.19
ChatGPT 0M - 11.62 11.29 82.39 37.48 62.05 5.28 38.63
ImgVAE - 72.06 12.58 12.05 79.95 42.38 63.55 1.52 6.34
Maria 135M 56.23 14.10 12.66 81.76 43.04 63.98 4.83 22.87
ZRIGF 658M 36.21 16.06 14.51 82.27 43.79 64.53 5.79 26.57
VIKDF 159M 14.98 16.47 14.84 82.54 43.84 64.88 6.53 35.12
Image-Chat Seq2Seq 24M 50.82 11.34 13.65 82.95 47.45 65.67 1.28 7.80
BART 406M 37.26 13.41 14.24 84.48 48.57 66.49 2.44 15.79
Llama-2 0M 193.20 9.93 11.56 85.44 40.18 63.05 4.69 30.81
ChatGPT 0M - 10.77 11.62 86.17 43.02 64.66 5.32 37.77
ImgVAE - 41.94 16.07 15.98 85.81 49.59 67.44 1.68 7.22
Maria 135M 37.49 14.74 14.59 85.72 50.58 66.89 2.57 11.99
[…]
ZRIGF 658M 29.82 16.86 17.21 86.30 51.41 68.56 2.59 10.62
[…]
VIKDF 159M 12.54 17.93 17.45 86.67 50.88 68.53 4.41 23.05
1/4 Data 159M 12.84 17.81 16.91 86.07 50.01 67.55 4.40 21.53
1/8 Data 159M 13.12 17.70 16.84 85.98 49.76 67.35 4.28 20.54
Zero Data 0M 27.32 15.35 15.30 85.04 45.86 66.02 6.17 32.93
4.3 Baseline Models
@@
  • 2.
    BART [45], a pre-eminent sequence-to-sequence pre-training model that utilizes Transformer architecture.
  • 3.
@@
Among them, ImgVAE, Maria, and ZRIGF are multimodal models, whereas the rest are unimodal text-based models. We have implemented all methods except for ImgVAE, where we directly used the outcomes from [4]. We prompt Llama-2 following the approach outlined in the Hugging Face blog [^1] https://huggingface.co/blog/llama2, and employ ChatGPT in accordance with the methodology described in [15].
4.4 Evaluation Metrics
@@ 5 Results and Discussion
5.1 Automatic Evaluation Results
Our proposed VIKDF demonstrates outstanding performance in zero-resource dialogue generation, outperforming both traditional and multimodal baseline models across various metrics on the Reddit Conversation and Image-Chat datasets. To ensure the robustness and reliability of our findings, we conducted additional experiments and statistical significance tests. Specifically, we ran VIKDF five times with different random seeds and performed one-sample t-tests comparing VIKDF to the best-performing baseline for each metric. Table 1 presents a comprehensive comparison of the automated evaluation metrics.
In the Reddit Conversation task, VIKDF achieves a significantly lower perplexity of 15.01, indicating superior fluency in generated dialogues compared to the nearest competitor, ZRIGF, which records a PPL of 36.21 in scenarios with explicit image guidance. Moreover, VIKDF outperforms all baselines in BLEU-1 and ROUGE-L scores, achieving 16.41 and 14.82, respectively, which highlights its ability to generate responses that are closely aligned with human references. Additionally, VIKDF surpasses other models in both Average and Greedy semantic similarity metrics, underscoring its enhanced ability to maintain semantic coherence in dialogues. Despite scoring slightly lower on the Extrema metric compared to ZRIGF, VIKDF remains competitive. Although it has a marginally lower Dis-2 score than ChatGPT, VIKDF’s strong performance in generating diverse dialogues is evident. This suggests that while maintaining high relevance and accuracy, VIKDF also generates a wide range of responses, contributing to more dynamic and engaging dialogues. Notably, VIKDF, with only 159M tuned parameters, significantly outperforms ZRIGF, which has 658M tuned parameters, in 7 out of 8 evaluation metrics. This indicates that VIKDF utilizes its parameters more efficiently, resulting in more coherent and contextually appropriate responses. Furthermore, VIKDF achieves superior performance compared to models that depend on explicit images, highlighting its effectiveness in enhancing dialogue capabilities without direct visual inputs. The outstanding performance of VIKDF in the Reddit Conversation task highlights the substantial benefits of incorporating implicit knowledge into purely text-based dialogue systems, proving that a deep fusion of visual and contextual understanding significantly enhances the quality and engagement of the dialogues generated.
@@ 5.1 Automatic Evaluation Results
Table 2: Human evaluation outcomes for the Image-Chat dataset in a fully zero-resource scenario.
@@
5.2 Human Evaluation Results